$SPAD/src/input gstbl.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool gstbl.output
)set message test on
)set message auto off
)clear all
--S 1 of 1
patrons: GeneralSparseTable(String, Integer, KeyedAccessFile(Integer), 0) := _
   table() ;
--R
--R
--R          Type: GeneralSparseTable(String,Integer,KeyedAccessFile(Integer),0)
--E 1
)spool
)lisp (bye)

———-

References^{[1] nothing}

4